OpenStudy (anonymous):

Please can someone explain me ho to rearrange the following formula to obtain t y =(V*sin(d))*t - 0.5*g*t^2 I trying to solve a problem by time and I can't see how to get this. here is the values in my work y = -35 d = 25 g = 9.8 v = 80 my book shows the final answer t = 7.814

OpenStudy (irishboy123):

\[y = V \sin(d)t - \frac{1}{2} g t^{2}\] \[g t^{2} -2 V \sin(d) t + 2y =0\] \[t = \frac{2V \sin(d) \pm \sqrt{4V^{2}\sin^{2}d - 4 (g)(2y)}}{2g}\] i suspect you have an easier way to do this eg use \[v^2 = u^2 + 2 a x \] as an intermediate step and then use \[v = u + at\]